# Online Process Reward Leanring for Agentic Reinforcement Learning

### v2

@@ Appendix A Algorithm
The algorithm flow of OPRL is detailed in Algorithm 1.
@@
Algorithm 1 Training LLM Agents with OPRL (GRPO as an example)
@@ Appendix B Environment Details
Appendix C Training Details
For hyperparameters, we use a constant learning rate $5\times 10^{-7}$ for policy models and $10^{-6}$ for PRMs with AdamW optimizer. Both policy and PRMs use a batch size of 64 and micro-batch size 8 for 200 training steps in veRL (Sheng et al., 2024). No KL-divergence penalty is applied. For WebShop and VisualSokoban, we apply a reward penalty of -0.1 to handle invalid actions generated by LLM agents. In WebShop and VisualSokoban, the maximum response length is 512 tokens, with the maximum prompt length 4096 tokens for WebShop and 1024 tokens for VisualSokoban. We sample 16 different groups per rollout in WebShop, resulting in a total of $16\times 8=128$ environments. In VisualSokoban, we sample 32 different groups per rollout, resulting in a total of $32\times 8=256$ environments. In SOTOPIA, the maximum prompt length is 6144 tokens and the maximum response length is 2048 tokens. All methods share the identical RL configurations for fair comparison, and all experiments are run on $8\timesA$100 GPUs.
@@ Appendix D Additional Results
Table 4 presents results of OPRL compared to baselines on WebShop using Qwen2.5-1.5B-Instruct as the base model.
Table 4: Performance on WebShop with Qwen2.5-1.5B-Instruct as the base model. Results are reported using one seed. OPRL still achieves the highest performance compared to baselines when transferring to small language models.

### v3

@@ Appendix A Algorithm
The algorithm flow of iStar is detailed in Algorithm 1.
@@
Algorithm 1 Training LLM Agents with iStar (GRPO as an example)
@@ Appendix B Environment Details
Appendix C Experiment Details
C.1 Training Details
In iStar, the implicit PRM is initialized from the base policy model, except for VisualSokoban, where the policy model uses Qwen2.5-VL-7B-Instruct and the PRM uses Qwen2.5-7B-Instruct. We use a constant learning rate $5\times 10^{-7}$ for the policy model and $10^{-6}$ for the implicit PRM with AdamW optimizer. Both the policy and the implicit PRM use a batch size of 64 and micro-batch size 8. No KL-divergence penalty is applied. We set the advantage coefficient $\alpha$ to 1 and $\beta=0.05$ for the implicit PRM training. The rollout size is set to 8 per prompt. All methods share the identical RL configurations for fair comparison, and all experiments are run on $8\timesA$100 GPUs. Below we show additional training details in each environment.
WebShop and VisualSokoban.
We use Qwen2.5-(VL)-7B-Instruct (Yang et al., 2024) as the base models for policy learning. To address invalid actions produced by LLM agents, a reward penalty of -0.1 is imposed. The maximum response length is 512 tokens, while the maximum prompt length is 4096 tokens in WebShop and 1024 tokens in VisualSokoban. We sample 16 different groups per rollout in WebShop, resulting in a total of $16\times 8=128$ environments. In VisualSokoban, we sample 32 different groups per rollout, resulting in a total of $32\times 8=256$ environments. Instead, PPO uses 128 and 256 separate environments for rollouts in WebShop and VisualSokoban, respectively. The rollout temperature is set to 1.0, while the validation temperature is set to 0.4. We implement experiments in veRL (Sheng et al., 2024), each for 200 training steps.
SOTOPIA.
We use Qwen2.5-7B-Instruct (Yang et al., 2024) and Llama3.1-8B-Instruct (Meta, 2024) as the base models for policy learning to demonstrate the robustness of our method to different model backbones. The maximum prompt length is 6144 tokens and the maximum response length is 2048 tokens. As with WebShop, we sample 16 different groups per rollout in WebShop, resulting in a total of $16\times 8=128$ environments (PPO uses 128 separate environments). The rollout temparature is set to 0.7. Each experiment implemented in veRL consists of 800 training steps.
C.2 Evaluation Prompts
We use ReAct (Yao et al., 2023) as the prompting strategy, with chain-of-thought (Wei et al., 2022) generated before each action. The prompt templates used for evaluating LLM or multimodal large language model (MLLM) agents in WebShop (Yao et al., 2022), VisualSokoban (Schrader, 2018), and SOTOPIA (Zhou et al., 2024) are presented in Figure 6, Figure 7, and Figure 8. Placeholders enclosed in curly braces({}) represent semantic slots, which are dynamically populated at runtime.
Figure 6: Evaluation prompts used for the LLM agent in WebShop.
Figure 7: Evaluation prompts used for the MLLM agent in VisualSokoban.
Figure 8: Evaluation prompts used for LLM agents in SOTOPIA.
@@ Appendix D Additional Results
Table 4 presents results of iStar compared to baselines on WebShop using Qwen2.5-1.5B-Instruct (Yang et al., 2024) as the base model.
Table 4: Performance on WebShop with Qwen2.5-1.5B-Instruct as the base model. Results are reported using one seed.